OpenStudy (anonymous):

Is the integral of 1/(sqrt(t) +sin(t)) from 0 to pi convergent?

OpenStudy (sirm3d):

\[\frac{1}{\sqrt{t}+\sin t}\leq \frac{1}{\sqrt{t}}\]

OpenStudy (anonymous):

Isn't \[t ^{-1/2}\] divergent?

OpenStudy (sirm3d):

yes. BUT the integral of \(t^{-1/2}\)?

OpenStudy (sirm3d):

don't forget that it is an integral problem...

OpenStudy (anonymous):

Ah, I get it now, Thanks again

OpenStudy (sirm3d):

yw.